OpenStudy (anonymous):

write the equation of the line that has a slope of -6 that passes through (7, 8).write the answer in slope-intercept form.

OpenStudy (saifoo.khan):

Use the formula: \[y- y_1 = m(x-x_1)\]Where m = slope. x1 = 7, y1 = 8.

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1\\ % (a,b) &({\color{red}{ 7}}\quad ,&{\color{blue}{ 8}})\quad \end{array} \\\quad \\ % slope = m slope = {\color{green}{ m}}= -6 \\ \quad \\ % point-slope intercept y-{\color{blue}{ y_1}}={\color{green}{ m}}(x-{\color{red}{ x_1}})\qquad \textit{plug in the values}\\ \qquad \uparrow\\ \textit{point-slope form}\)

OpenStudy (bee_see):

or you can do./